OpenStudy (samirahdanyel):

A Ferris wheel has a diameter of 42 feet. It rotates 3 times per minute. Approximately how far will a passenger travel during a 5-minute ride? 132 feet 659 feet 1,978 feet 3,956 feet

OpenStudy (zzr0ck3r):

so circle with radius 21 has circumference is 2pi*r and that is how much we will travel in one turn, we go 3 turns per min for five mins we have 15 turns. can you get it from here?

OpenStudy (samirahdanyel):

no I don't

OpenStudy (samirahdanyel):

oh wait I get it

OpenStudy (samirahdanyel):

Now i gotta figure it out for 5 min ride?

OpenStudy (samirahdanyel):

which I dont know how to do

jimthompson5910 (jim_thompson5910):

@samirahdanyel what did you get for the circumference of the circle?

OpenStudy (samirahdanyel):

I didn't do that yet.

jimthompson5910 (jim_thompson5910):

use 2*pi*r pi = 3.14 (approx) r = 21 since the diameter is 42 and 42/2 = 21

OpenStudy (samirahdanyel):

I got 42 pi

jimthompson5910 (jim_thompson5910):

so one full cycle has the person traveling `42pi` feet in one minute the wheel rotates 3 times, so how many feet does this person go in one minute?

OpenStudy (samirahdanyel):

Do i multiply

OpenStudy (samirahdanyel):

or something like that

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

3*42pi = ??

jimthompson5910 (jim_thompson5910):

or you can think of it as 42pi+42pi+42pi but that is a bit more cumbersome

OpenStudy (samirahdanyel):

I got 126 pi

jimthompson5910 (jim_thompson5910):

yep, so in one minute, the person travels `126pi` feet

jimthompson5910 (jim_thompson5910):

now multiply that by 5 to get the total distance they travel in 5 minutes

OpenStudy (samirahdanyel):

i got 630 pi

jimthompson5910 (jim_thompson5910):

me too, so that is the answer in terms of pi

jimthompson5910 (jim_thompson5910):

630pi means 630*pi replace pi with 3.14

OpenStudy (samirahdanyel):

okay so i do 630 (3.14) *5

jimthompson5910 (jim_thompson5910):

no just 630*3.14

jimthompson5910 (jim_thompson5910):

the 5 is already built in

OpenStudy (samirahdanyel):

okay

OpenStudy (samirahdanyel):

I got 1979.20337

jimthompson5910 (jim_thompson5910):

If you use pi = 3.14, then you should get 630*pi = 630*3.14 = 1,978.2

OpenStudy (samirahdanyel):

oh i didnt put it in parantheses